OpenStudy (anonymous):

ok here's my issue, if the building is 285 ft tall, and free falling objects neglecting air resistance, fall about 10m/s. Dividing that out, it comes to about 9.5 seconds. Am I doing something wrong with the equation because if t=2d/g^1/2 i got about 28.5 seconds. So which is the right answer?

OpenStudy (amistre64):

when dealing with feet; it tends to be best to use the -16t^2 for footric measures as opposed to metric stuff

OpenStudy (amistre64):

-16t^2 + initial speed (t) + initial height is the usual formula

OpenStudy (anonymous):

I'm so confused, to be honest I am not that smart, it prolly woulda been easier if I had taken calculus first but this is new to me considering I'm still in elementary algebra to bring up to college standard. I don't know what what all that means

OpenStudy (amistre64):

first off, what is it your asking above? just so that i know if im on the right track

OpenStudy (anonymous):

ok if I dropped a rock of a building 285 ft tall how long would it take to hit the ground?

OpenStudy (amistre64):

thats doable :) there is an equation that the smart guys have already developed and it goes like this: effect of gravity x time^2 plus, starting speed of the object x time plus, starting height of object

OpenStudy (amistre64):

the effect of gravity is in feet here so we use -16 t^2, where t is where we put in our time. the initial speed of your rock is zero so that part goes away and were are left with the part that says how high it started from -16t^2 +285

OpenStudy (amistre64):

now when that equation for the height of the rock = 0; we have hit the ground

OpenStudy (anonymous):

I know, I just feel stupid I know the equations they're just not turning out with a reasonable answer

OpenStudy (amistre64):

-16t^2 +285 = 0 ; subtract 285 from both sides -285 -285 ------------------ -16t^2 = -285 ; divide off the -16 /-16 /-16 ------------------ t^2 = 285/16 ; now take the sqaure root of each side

OpenStudy (amistre64):

t = sqrt(285)/4 simplify as wanted

OpenStudy (anonymous):

ok so it's 4.21 seconds?

OpenStudy (anonymous):

or did I do that wrong?

OpenStudy (amistre64):

its about 4.21 secs yes

OpenStudy (amistre64):

4.22 maybe with the rounding

OpenStudy (anonymous):

ok so now I'm more confused. in the book it says free falling objects fall at 10 m/s andthere are 95 meters in 285 ft. Wouldn't it be closer to 10 seconds?

OpenStudy (amistre64):

-4.9t^2 + 95 ; assuming your feet to meters conversion is good t = sqrt(95/4.9) =abt 4.40315 .... get a more exact figure on the metric but otherwise they are about the same

OpenStudy (amistre64):

even if its 10; half that is 5 and sqrt(95/5) is still abt 4.5 sec

OpenStudy (anonymous):

ok, well thank you so much for your help. I really appreciate it

OpenStudy (amistre64):

285 = 86.87 meters to be more exact. all the errors are from aproximating .. but they tend to form a concensus of about 4.5 seconds